OpenStudy (anonymous):

Part 1: Given log a b = x, replace the variables a and b with integers of your choice. (2 points) Part 2: Show your work to solve the logarithm you created using the change of base formula. (4 points) Part 3: Provide the solution. (2 points)

hero (hero):

1. If a = 3, and b = 9, then \[\log_a(b) = x\] becomes: \[\log_3(9)= x\] 2. If you use the change of base formula, then \[\log_3(9)= x\] becomes \[3^x = 9\] 3. Since \[3^2 = 9\] x = 2

OpenStudy (anonymous):

But is that the change of base formula? Dont you have to make it into a fraction?

hero (hero):

1. If a = 3, b = 9, then \[\log_3(9) = x\] 2. Using change of base formula: \[\log_3(9) = x\] becomes: \[\frac{\log 9}{\log 3} = x\] 3. Using your calculator, you'll get x = 2

hero (hero):

I consider either one of them the change of base formula.

OpenStudy (anonymous):

Thank You SO MUCH!!!! I owe you:)

hero (hero):

Now I just showed you two ways of doing the same problem. You should have no trouble with doing any similar problems.

OpenStudy (anonymous):

:)